# Yaroslav Aksenov, Ian Maksimov, Nikita Balagansky, Daniil Gavrilov Tinkoff

### v2

@@ 5 Discussion
This paper introduces a new class of language model alignment methods called Trust Region (TR). TR-DPO, TR-IPO and TR-KTO stand out from classical offline optimization methods by updating the reference policy during the training process. The "Trust Region" designation reflects our approach’s foundational principle of allowing models to explore beyond the initial SFT policy to achieve improved alignment convergence. The effectiveness of this strategy is likely based on the nuanced dynamics of the loss function’s gradient behavior. We validate the proposed technique on a variety of tasks and model sizes. The results show that our modifications improve performance in the majority of scenarios. This underscores the universality and reliability of our approach.
Limitations and future work. In our study, we propose a new paradigm for language model alignment and demonstrate its effectiveness on two of the most popular datasets. Future research could explore how the proposed method generalizes to other domains, modalities, and dataset sizes, especially small datasets. In this work, we show that models trained by our method suffer less from an increase in response length, but the justification of this behavior is beyond the scope of our work and requires further investigation. Finally, our results are based on an automatic evaluation by GPT-4. This raises the question of how suitable such evaluation is for assessing alignment techniques, and whether there could be more suitable evaluation methods.
Societal impacts. The methods proposed in this article, like other alignment methods, can be used to create unsafe assistants. We are only warning about this possibility, but our research is aimed at creating helpful and harmless assistants, and we expect that our methods will be used for the same purposes.
Appendix A Training Details
@@
The training of Pythia models adhered to a set of hyperparameters optimized for performance (see Table 2). Unless otherwise noted, the following hyperparameters were consistent across all training setups.
@@
Max Tokens Length 1024
@@
Optimizer Adam (Kingma and Ba,, 2014)
@@
Learning Schedule Linear Decay (Loshchilov and Hutter,, 2016)
@@
Memory optimization DeepSpeed (Rasley et al.,, 2020)
Attention Mechanism Flash Attention 2 (Dao,, 2023)
Table 2: Training hyperparameters for Pythia Models
All computations were performed on 8x NVIDIA A100 GPUs with 80GB of memory, which provided the necessary computational power to efficiently train our models. Depending on the number of parameters of pre-trained models, training took from $3$ to $12$ hours for each model.
@@
Appendix B License
The Anthropic HH and Reddit TL;DR datasets utilized during the course of our study are licensed under the terms of the Massachusetts Institute of Technology (MIT) License. You can learn more about this license at https://opensource.org/licenses/MIT.
@@ Appendix B License
Appendix C Human-centric analysis
Figure 7: Scores of standardized HC metrics (see Section 4.1) across a range of $\tau$ values at $2^{n}$ intervals, where $n=6,\ldots,10$, in TR-DPOτ. The analysis demonstrates that $\tau$ values of 512 and 1024 consistently outperform the DPO baseline, as evidenced by higher bars representing superior performance across HC metrics.
C.1 TR-IPO Analysis
@@
Figure 8: Scores of standardized HC metrics (see Section 4.1) across a range of $\alpha$ values [0.1; 0.8] in TR-IPOα (top) and $\tau$ values at $2^{n}$ intervals, where $n=6,\ldots,10$, in TR-IPOτ (top). The analysis demonstrates that $\alpha$ values between $0.5$ and $0.7$, and $\tau$ values of $512$ and $1024$ consistently outperform the IPO baseline, as evidenced by higher bars representing superior performance across HC metrics.
@@
Figure 9: Based on PoI analysis, TR-IPOα and TR-IPOτ outperform IPO across HC metrics such as coherence, correctness, helpfulness, and harmlessness. For $\alpha=0.5$, $0.6$, $0.7$ and $\tau=256$, $512$ and $1024$, the confidence intervals do not cross the 0.5 probability line, denoting statistical significance of the enhancements. TR-DPOα spans $\alpha$ values [0.1; 0.8]; TR-DPOτ tests $\tau$ at $2^{n}$ intervals, $n=6,\ldots,10$ with the Pythia 2.8B model. See Section 4.3 for more details.
C.2 TR-KTO Analysis
The $\alpha$ value of 0.6 is significantly better than the KTO baseline for the TR-KTO method, as evidenced by the PoI analysis in Figure 11(a), with confidence intervals that do not cross the 0.5 probability threshold. For $\tau$ settings, a value of 512 yielded the best results, depicted in Figure 11(b). This setting significantly improved coherence, correctness, helpfulness and harmlessness, as shown in Figure 10
Figure 10: Scores of standardized HC metrics (see Section 4.1) across a range of $\alpha$ values [0.1; 0.8] in TR-KTOα (top) and $\tau$ values at $2^{n}$ intervals, where $n=7,\ldots,10$, in TR-KTOτ (bottom). The analysis demonstrates that the $\alpha$ value of $0.6$ and the $\tau$ value of $512$ outperform the KTO baseline, as evidenced by higher bars representing superior performance across coherence, correctness, and harmlessness metrics.
@@ C.2 TR-KTO Analysis
Appendix D Model Dynamics and Policy Divergence
D.1 TR-IPO Analysis
The dependency of the average HC metrics and the KL divergence on the $\beta$ parameter for the IPO, TR-IPOα, and TR-IPOτ methods is illustrated in Figure 12(a) using the Pythia 2.8B model on the Anthropic-HH dataset. It is observed that the peak value of HC metrics for the IPO method is attained at $\beta=0.01$, which is the setting used in the experiments described in Sections 4.2 and 4.3. Note that the HC metrics for the IPO method decline more rapidly after this peak compared to its modified versions, allowing the TR methods to yield superior results.
Figure 12: (a) The relationship between KL divergence and HC mean value for IPO and TR-IPO ($\alpha=0.6$, $\tau=512$) across different $\beta$ values. The best HC value for IPO occurs at $\beta=0.01$ but declines beyond this point. TR-IPO methods extend further in KL divergence without a rapid HC decline. (b) Gradient scale variations for IPO, TR-IPO with a soft update ($\alpha=0.6$), and TR-IPO with a hard update ($\tau=512$) at $\beta=0.01$. The soft update achieves a gradual increase, while the hard update acts as a gradient reset. See Section 4.4 for details.
The gradient of the IPO loss function, as defined in Equation 3, is similar to the DPO gradient (Equation 8). It can be represented with a weighting factor preceding the gradient term in Equation 9 as follows:
$\nabla_{\theta}\mathcal{L}_{\text{IPO}}(\pi_{\theta},\pi_{\text{ref}})=\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[2\left(\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\frac{1}{2\beta}\right)\nabla_{\theta}\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\theta}(y_{l}|x)}\right].$ (9)
The absolute value of the term $2\left(\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\frac{1}{2\beta}\right)$ acts as a weighting factor, shifting the training policy to maintain a margin equal to $\frac{1}{2\beta}$ in terms of the log probabilities for chosen and rejected samples. This mechanism effectively adjusts the influence of each training sample based on how the current policy $\pi_{\theta}$ assigns probabilities to outcomes compared to the reference model $\pi_{\text{ref}}$. When $\pi_{\theta}$ significantly favors the preferred outcome $y_{w}$ over what $\pi_{\text{ref}}$ does, or underestimates $y_{l}$, the term grows larger, promoting larger updates to reinforce or adjust these disparities. Conversely, smaller disparities result in smaller updates.
Similar to the DPO dynamics (Figure 6(b)), the smoothed gradient scales are depicted in Figure 12(b). We hypothesize that there is a relationship between these observed gradient scales and the training dynamics.
D.2 TR-KTO Analysis
The relationship between the average HC metrics and the KL divergence with different $\beta$ parameters is explored for the KTO, TR-KTOα, and TR-KTOτ methods in Figure 13(a), utilizing the Pythia 2.8B model on the Anthropic-HH dataset. It is observed that the HC metrics for the KTO method peak at $\beta=0.05$, the setting used in the experiments described in Sections 4.2 and 4.3. Beyond this peak, the HC metrics for the KTO method decline more rapidly than those for its modified versions, thereby allowing the TR methods to achieve superior outcomes.
Figure 13: (a) The relationship between KL divergence and HC mean value for KTO and TR-KTO ($\alpha=0.6$, $\tau=512$) across different $\beta$ values. The best HC value for KTO occurs at $\beta=0.05$ but declines beyond this point. TR-KTO methods maintain further KL divergence without a rapid decline in HC. Figures (b) and (c) display the gradient dynamics for chosen and rejected scenarios respectively for KTO, TR-KTO with a soft update ($\alpha=0.6$), and TR-KTO with a hard update ($\tau=512$) at $\beta=0.05$. The soft update results in a gradual increase, while the hard update serves as a gradient reset for both chosen and rejected terms. See Section 4.4 for details.
The gradient function for the KTO model, as defined in Equation 10, introduces a distinct approach to handling the chosen and rejected sample outputs, reflected through separate gradient terms. This differentiation is necessary due to the structure of the KTO loss function (Equation 4), which explicitly addresses the decision-making biases outlined by Kahneman and Tversky, (1979). The KTO gradient comprises two key components, each addressing a different aspect of policy behavior:
$\begin{split}\nabla_{\theta}\mathcal{L}_{\text{KTO}}(\pi_{\theta},\pi_{\text{ref}})&=-\beta\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}[\\ &\quad\lambda_{w}\sigma(z(y_{w}))\sigma(-z(y_{w}))\nabla_{\theta}\log\pi_{\theta}(y_{w}|x)\\ &\quad-\lambda_{l}\sigma(z(y_{l}))\sigma(-z(y_{l}))\nabla_{\theta}\log\pi_{\theta}(y_{l}|x)],\end{split}$ (10)
where $z(y)=\beta\log\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}-z_{\text{ref}}$, and $z_{\text{ref}}=\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(\cdot|x)}\big{[}\beta\log\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}\big{]}$, $\lambda_{w}$ and $\lambda_{l}$ are coefficients controlling the degree of loss aversion. According to Ethayarajh et al., (2024), the coefficients should possess the following property: $\frac{\lambda_{w}}{\lambda_{l}}\in\Big{[}1,\frac{4}{3}\Big{]}$. Therefore, for our experiments, we take coefficients equal to 1.
Figures 13(b) and 13(c) show the absolute values of gradient terms for chosen and rejected terms according to Equation 10. Unlike the DPO and IPO gradient functions (Equations 8 and 9), KTO requires distinct visualization of each gradient component’s influences on the overall optimization process.
D.3 Impact of Update Strategies on Gradient Dynamics in TR-DPO
Informed by Equation 8, we analyze how the magnitude of gradients behaves under different TR-DPO update strategies. Our observations, illustrated in Figures 14(a) and 14(b), reveal that higher $\alpha$ values in soft updates and lower $\tau$ values in hard updates enhance the gradient scales when compared to the DPO baseline. This pattern suggests that updates closer to the baseline policy tend to stabilize the training process, while more aggressive updates lead to increased variability in gradient magnitude. Soft updates incrementally escalate the gradients, and hard updates introduce distinct spikes at each update step.

### v3

@@ 5 Discussion
This paper introduces Trust Region (TR) methods—TR-DPO, TR-IPO, and TR-KTO—that update the reference policy during training, distinguishing them from classical offline optimization methods. Our approach addresses the reward overoptimization problem, showing that TR methods are less affected by overoptimization compared to their vanilla counterparts.
Limitations and Future Work. Future research could explore the generalization of our methods to other domains, modalities, and especially smaller datasets. In particular, large $\tau$ values in hard updates may not scale well, indicating the need for more flexible, data-adaptive approaches.
We linked reward overoptimization to gradient dynamics that decrease in-domain sequence probabilities and demonstrated that our approach improves offline alignment methods. Other methods to prevent overoptimization (e.g., scheduling $\beta$ values or optimizing chosen and rejected sequences with different weights) could also be explored.
Our results rely on automatic evaluation by GPT-4, raising questions about the suitability of such methods for assessing alignment techniques and whether more appropriate evaluation methods exist.
Appendix A Derivation of DPO Hessian
In this section, we derive the Hessian of the DPO (Direct Preference Optimization) loss. We start with the gradient of the DPO loss:
$\begin{split}&\nabla_{\theta}\mathcal{L}_{\text{DPO}}(\pi_{\theta},\pi_{\text{ref}})=\\ &=-\beta\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\sigma\left(\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\theta}(y_{w}|x)}-\beta\log\frac{\pi_{\text{ref}}(y_{l}|x)}{\pi_{\text{ref}}(y_{w}|x)}\right)\nabla_{\theta}\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\theta}(y_{l}|x)}\right].\end{split}$ (10)
Let us denote $s=\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\theta}(y_{w}|x)}-\beta\log\frac{\pi_{\text{ref}}(y_{l}|x)}{\pi_{\text{ref}}(y_{w}|x)}$, and $g=\nabla_{\theta}\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\theta}(y_{l}|x)}$. Note that $g=-\frac{1}{\beta}\nabla_{\theta}s$. Therefore,
$\begin{split}\nabla^{2}_{\theta}\mathcal{L}_{\text{DPO}}(\pi_{\theta},\pi_{\text{ref}})=-\beta&\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\sigma(s)\nabla_{\theta}g+g\sigma(s)\big{(}1-\sigma(s)\big{)}\big{(}\nabla_{\theta}s\big{)}^{\top}\right]=\\ =-\beta&\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\sigma(s)(-\frac{1}{\beta}\nabla^{2}_{\theta}s)+(-\frac{1}{\beta}\nabla_{\theta}s)\sigma(s)\big{(}1-\sigma(s)\big{)}\big{(}\nabla_{\theta}s\big{)}^{\top}\right]=\\ =&\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\sigma(s)\nabla^{2}_{\theta}s+\sigma(s)\big{(}1-\sigma(s)\big{)}\nabla_{\theta}s\big{(}\nabla_{\theta}s\big{)}^{\top}\right].\end{split}$ (11)
Appendix B Toy Example Details
In real-world scenarios, the size of the token vocabulary and the number of parameters in a model are quite large. To simplify our analysis, we utilized the setup suggested by Rafailov et al. (2024a). Essentially, we used a language model to simulate a Markov Decision Process (MDP), as visualized in Figure 7. In each state, there are three potential actions, each leading to a new state, except for those at depth 2 (with the initial state at depth 0) — these all lead to a single terminal state.
Figure 7: The MDP used in the toy example to analyze the overoptimization problem. In all states, there are three possible actions that lead to new states. Leaf nodes lead to a single terminal state.
We used a Recurrent Neural Network (RNN) with a hidden size of 3 and an output size of 3 to predict the actions. The input was a sequence of actions encoded as one-hot vectors of dimension 4 (representing the 3 actions and a "beginning of sequence" token).
The RNN training comprised two classical stages: Supervised Fine-Tuning (SFT) and offline alignment. To simulate a real situation where the number of texts in the datasets is considerably less than the set of possible responses, we used a set of three trajectories for the SFT stage: $(a_{0},a_{0},a_{0})$, $(a_{1},a_{1},a_{0})$, and $(a_{2},a_{2},a_{2})$. For the alignment stage, we used a set consisting of only one pair: $(a_{1},a_{1},a_{0})\succ(a_{0},a_{0},a_{0})$. Both trajectories in the pair are included in the SFT dataset. For soft updates, we used $\alpha=0.6$, and for hard updates, we used $\tau=8$.
(a) IPO
[…]
(a) KTO
[…]
Appendix C Probability Analysis on the Real Task
Similarly to the toy example outlined in Appendix B, we opted to explore the distribution of probability mass between in-domain and out-of-domain sequences. Since it is unfeasible to visualize the entire OOD text space in real-world scenarios, we inspect only the behavior of probabilities for chosen and rejected texts during the training of Pythia 2.8B using both DPO and TR-DPO methods on the Anthropic-HH dataset.
The selected hyperparameters were $\alpha=0.6$ and $\tau=512$ due to their optimal performance on this task (see Section 4.2). According to Rafailov et al. (2024a), as the KL divergence increases with the SFT policy, there is a decrease in the log probability of both chosen and rejected texts (refer to Section 3.1). This necessitates the selection of $\beta$ values for comparison such that the methods reach an equivalent KL divergence with the SFT policy. Following Figure 5, $\beta$ was chosen as 0.03 for TR-DPO and 0.01 for DPO. At these $\beta$ values, there is a significant difference in HC metrics (indicating overoptimization in DPO), while maintaining the same KL divergence with the SFT policy.
As depicted in the plots, probabilities for both chosen and rejected texts are higher for TR methods, indicating less redistribution of probability mass onto OOD texts. This could suggest a reduced tendency to overoptimization.
It is noteworthy that even though the probabilities of rejected texts remain higher for TR methods, these methods demonstrate better performance, as shown in Appendix Table 8. An interesting conclusion from this fact is that for optimal model performance, it is more crucial to prevent probability leakage to OOD examples during training than to minimize the probability of rejected texts.
(a) Chosen
(b) Rejected
Figure 10: The batch-averaged logarithm values of probabilities for chosen (a) and rejected (b) texts from the Anthropic-HH dataset. The vanilla DPO method is compared to its TR modifications with hyperparameters $\alpha=0.6$ and $\tau=512$. To ensure the methods were equally distant from the SFT policy according to KL divergence, $\beta$ values of 0.03 for TR methods and 0.01 for DPO were used. The probability values for chosen and rejected texts in TR-DPO are higher than those in DPO, suggesting that less probability mass has transitioned to OOD texts. This result indicates that our proposed methods have a reduced tendency towards overoptimizing. For more details, refer to Appendix C.
Appendix D Implementation Details
D.1 Training Details
@@
The training of Pythia and Llama models adhered to a set of hyperparameters optimized for performance (see Table 3). Unless otherwise noted, the following hyperparameters were consistent across all training setups.
@@
Max Tokens Length 1024 (Pythia), 2048 (Llama3)
@@
Optimizer Adam (Kingma & Ba, 2014)
@@
Learning Schedule Linear Decay (Loshchilov & Hutter, 2016)
@@
Memory optimization DeepSpeed (Rasley et al., 2020)
Attention Mechanism Flash Attention 2 (Dao, 2023)
Table 3: Training hyperparameters for Pythia and Llama3 Models
All computations were performed on 8 NVIDIA A100 GPUs with 80GB of memory, which provided the necessary computational power to efficiently train our models. Depending on the number of parameters of pre-trained models, training took between 3 to 12 hours for each model.
@@
A summary of the dataset sizes is provided in Table 4. The Reddit TL;DR dataset was processed to remove duplicates, retaining only uniquely preferred summaries for SFT.
Dataset Training Examples Validation Examples
Anthropic-HH 160,800 8,552
Reddit TL;DR summarization (SFT) 41,947 11,941
Reddit TL;DR summarization (Preference) 73,396 21,198
UltraChat-200k 207,865 23,110
UltraFeedback 61,135 2,000
Table 4: Summary of dataset sizes used for training and validation
D.2 Generation Details
For the generation of outputs, both Pythia and Llama models utilized optimized generation hyperparameters to ensure coherent and high-quality text. The settings for each model are outlined below.
Hyperparameter Pythia Models Llama3 Models
Temperature 1.0 0.9
Top-k 40 40
Top-p 0.7 1.0
Max New Tokens 512 2048
Table 5: Generation hyperparameters for Pythia and Llama3 Models
Appendix E Ablation study on TR Hyperparameters
The ablation study demonstrates that the performance of TR methods is sensitive to the choice of hyperparameters $\alpha$ and $\tau$, with varying optimal values depending on the task and model size. For the Helpful&Harmless task on the Pythia 2.8B model, $\alpha=0.6$ and $\tau=512$ provided consistent improvements over baseline methods (see Table 6), affirming their near-optimality for this setup. However, for larger models and tasks such as AlpacaEval 2, the optimal hyperparameters shifted, with $\alpha=0.7$ and $\alpha=0.8$ performing better on the Llama3 models, and $\tau=32$ or $\tau=64$ yielding improved results (see Table 7). Notably, $\alpha=0.9$ and $\tau$ values smaller than 16 caused instability in both training and generation processes, leading to their exclusion from the tables presented.
These results suggest that while $\alpha=0.6$ and $\tau=512$ provide a reasonable baseline for general tasks, tuning hyperparameters for specific tasks and model configurations can further enhance performance. The sensitivity of TR methods to hyperparameters should be viewed as an opportunity for fine-tuning rather than a limitation, as the proposed methods consistently outperform baseline counterparts across a range of configurations.
TR-DPOα vs. DPO TR-IPOα vs. IPO TR-KTOα vs. KTO
$\alpha$ Win Tie Lose Win Tie Lose Win Tie Lose
0.1 40.0 22.2 37.8 40.0 24.8 35.2 11.0 10.4 78.6
0.2 38.2 25.0 36.8 42.8 23.0 34.2 14.0 10.2 75.8
0.3 38.2 25.4 36.4 42.4 24.2 33.4 13.8 10.4 75.8
0.4 41.4 22.6 36.0 46.2 21.0 32.8 35.8 29.0 35.2
0.5 42.8 22.8 35.4 45.6 22.8 31.6 35.8 27.2 37.0
0.6 42.4 25.2 32.4 43.2 22.4 34.4 37.4 28.2 34.4
0.7 42.2 22.2 35.6 46.4 21.8 31.8 14.0 10.8 75.2
0.8 39.2 19.2 41.6 34.2 18.0 47.8 6.0 8.0 86.0
TR-DPOτ vs. DPO TR-IPOτ vs. IPO TR-KTOτ vs. KTO
$\tau$ Win Tie Lose Win Tie Lose Win Tie Lose
64 29.6 19.4 51.0 31.2 15.8 53.0 2.6 9.4 88.0
128 39.8 20.2 40.0 46.8 18.0 35.2 30.8 19.0 50.2
256 41.0 24.0 35.0 44.2 21.8 34.0 33.2 29.2 37.6
512 41.8 24.0 34.2 45.4 22.4 32.2 40.4 26.0 33.6
1024 42.8 22.6 34.6 43.4 20.8 35.8 11.8 9.2 79.0
Table 6: Ablation study on the hyperparameters of the TR methods, through AutoSxS comparison with classical counterparts, utilizing Pythia 2.8B on the Anthropic-HH dataset.
TR-DPOα TR-IPOα TR-KTOα
$\alpha$ LC(std) WR(std) LC(std) WR(std) LC(std) WR(std)
Base 15.86 (0.77) 11.23 (0.96) 15.87 (0.74) 11.49 (0.97) 15.81 (0.75) 12.99 (1.02)
0.1 17.23 (0.79) 12.17 (1.01) 15.11 (0.74) 11.15 (0.99) 18.01 (0.82) 14.33 (1.09)
0.2 16.80 (0.76) 12.82 (1.03) 16.86 (0.76) 12.71 (1.01) 16.52 (0.79) 13.25 (1.02)
0.3 19.83 (0.82) 14.89 (1.10) 20.20 (0.85) 15.39 (1.11) 18.15 (0.80) 14.37 (1.06)
0.4 19.40 (0.81) 15.01 (1.11) 18.85 (0.79) 14.56 (1.06) 19.73 (0.82) 16.18 (1.12)
0.5 21.58 (0.80) 18.41 (1.14) 21.51 (0.80) 18.56 (1.15) 18.19 (0.73) 14.99 (1.06)
0.6 22.95 (0.80) 19.54 (1.17) 22.41 (0.81) 19.30 (1.18) 20.70 (0.83) 17.31 (1.12)
0.7 24.32 (0.83) 19.37 (1.18) 25.26 (0.86) 21.18 (1.25) 19.61 (0.80) 15.68 (1.09)
0.8 27.25 (0.85) 23.92 (1.25) 29.48 (0.84) 25.36 (1.29) 14.66 (0.70) 12.54 (0.99)
TR-DPOτ TR-IPOτ TR-KTOτ
$\tau$ LC(std) WR(std) LC(std) WR(std) LC(std) WR(std)
Base 15.86 (0.77) 11.23 (0.96) 15.87 (0.74) 11.49 (0.97) 15.81 (0.75) 12.99 (1.02)
16 26.36 (0.75) 23.22 (1.27) 25.11 (0.73) 22.15 (1.23) 11.78 (0.55) 11.41 (0.99)
32 27.71 (0.84) 25.73 (1.28) 28.08 (0.83) 25.42 (1.29) 16.54 (0.71) 13.85 (1.05)
64 24.30 (0.79) 21.98 (1.24) 25.40 (0.78) 22.84 (1.24) 20.34 (0.81) 16.59 (1.13)
128 21.70 (0.79) 18.55 (1.16) 21.56 (0.76) 18.07 (1.15) 20.22 (0.82) 16.17 (1.11)
256 20.25 (0.81) 16.36 (1.11) 19.34 (0.78) 16.27 (1.10) 20.70 (0.83) 18.00 (1.14)
Table 7: Benchmark results from Alpaca Eval, comparing various hyperparameters of TR methods on the Llama3 models, utilizing UltraFeedback and UltraChat datasets.
Appendix F License
The Anthropic-HH, Reddit TL;DR, UltraChat 200k, and UltraFeedback datasets utilized during the course of our study are licensed under the terms of the Massachusetts Institute of Technology (MIT) License. You can learn more about this license at https://opensource.org/licenses/MIT.
@@ Appendix F License
The Llama3 models used as pretrained models in the study are licensed under the Meta Llama 3 Community License Agreement. For more details on this license, you can refer to https://www.llama.com/llama3/license/.
Appendix G Human-centric Analysis
G.1 TR-DPO Analysis
Figure 12: Standardized HC metrics (see Section 4.1) across a range of $\tau$ values at $2^{n}$ intervals, where $n=6,\ldots,10$, in TR-DPOτ. The analysis demonstrates that $\tau$ values of 512 and 1024 consistently outperform the DPO baseline, as evidenced by higher bars representing superior performance across HC metrics.
G.2 TR-IPO Analysis
@@
Figure 13: Standardized HC metrics (see Section 4.1) across a range of $\alpha$ values [0.1; 0.8] in TR-IPOα (top) and $\tau$ values at $2^{n}$ intervals, where $n=6,\ldots,10$, in TR-IPOτ. The analysis demonstrates that $\alpha$ values between $0.5$ and $0.7$, and $\tau$ values of $512$ and $1024$ consistently outperform the IPO baseline, as evidenced by higher bars representing superior performance across HC metrics.
@@
Figure 14: Based on PoI analysis, TR-IPOα and TR-IPOτ outperform IPO across HC metrics such as coherence, correctness, helpfulness, and harmlessness. For $\alpha=0.5$, $0.6$, $0.7$ and $\tau=256$, $512$, and $1024$, the confidence intervals do not cross the 0.5 probability line, denoting statistical significance of the enhancements. TR-DPOα spans $\alpha$ values [0.1; 0.8]; TR-DPOτ tests $\tau$ at $2^{n}$ intervals, $n=6,\ldots,10$ with the Pythia 2.8B model. See Section 4.4 for more details.
G.3 TR-KTO Analysis
The $\alpha$ value of 0.6 is significantly better than the KTO baseline for the TR-KTO method, as evidenced by the PoI analysis in Figure 16(a), with confidence intervals that do not cross the 0.5 probability threshold. For $\tau$ settings, a value of 512 yielded the best results, depicted in Figure 16(b). This setting significantly improved coherence, correctness, helpfulness, and harmlessness, as shown in Figure 15.
Figure 15: Standardized HC metrics (see Section 4.1) across a range of $\alpha$ values [0.1; 0.8] in TR-KTOα (top) and $\tau$ values at $2^{n}$ intervals, where $n=7,\ldots,10$, in TR-KTOτ. The analysis demonstrates that the $\alpha$ value of $0.6$ and the $\tau$ value of $512$ outperform the KTO baseline, as evidenced by higher bars representing superior performance across coherence, correctness, and harmlessness metrics.
@@ G.3 TR-KTO Analysis
Appendix H Impact of Update Strategies on Gradient Dynamics in TR-DPO
Informed by Equation 6, we analyze how the coefficient $\sigma\left(\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\theta}(y_{w}|x)}-\beta\log\frac{\pi_{\text{ref}}(y_{l}|x)}{\pi_{\text{ref}}(y_{w}|x)}\right)$ behaves under different TR-DPO update strategies. Our observations, illustrated in Figures 17(a) and 17(b), reveal that higher $\alpha$ values in soft updates and lower $\tau$ values in hard updates enhance the gradient scales compared to the DPO baseline. The behavior of the coefficients in these plots confirms the idea presented in Section 3.1 that updating the reference policy allows us to "reset" the optimization process, increasing not only the scale of the loss function gradient but, more importantly, its Hessian.